OpenStudy (anonymous):

The local movie theater has a special movie night on Wednesdays. Admission is $4.00 for any age moviegoer. Concessions are $2 per item. Which of the following equations best represents the relationship between the total cost, y, and the number of concessions bought, x. Medals to any people with good answers!! :)

OpenStudy (anonymous):

Oh the choices are: a. y = 4x + 2 b. x = 4 + 2y c. y = 2x + 4

OpenStudy (anonymous):

c

OpenStudy (eyust707):

would you like an explanation why?

OpenStudy (anonymous):

You don't have to if you don't want to. :)

OpenStudy (anonymous):

because x (number of concessions) changes

OpenStudy (anonymous):

and it should always be set equal to y

OpenStudy (eyust707):

lol

OpenStudy (anonymous):

B. 4 is the admission cost, and 2Y shows that per Y, there are 2 dollars.

OpenStudy (anonymous):

C then*

OpenStudy (anonymous):

So it's not c then?

OpenStudy (anonymous):

and because the admission cost doesnt change

OpenStudy (eyust707):

the correct answer is c

OpenStudy (anonymous):

It's C.